The ideal diameter of a steel reinforcement rod is 2.8cm. The actual diameter may vary from the ideal by at most 0.006cm. Find the range of acceptable diameters for the steel rod.
Answer by Fombitz(32388)

Since the diameter can vary from the nominal value of 2.8cm both in the negative direction and in the positive direction, find the two limits.
Upper limit = 2.800+0.006 = 2.806 cm
Lower limit = 2.800-0.006 = 2.794 cm
Acceptable diameter range is [2.794-2.806] cm.